# Balance Reward and Safety Optimization for Safe Reinforcement Learning: A Perspective of Gradient Manipulation

### v1

@@ Abstract
Ensuring the safety of Reinforcement Learning (RL) is crucial for its deployment in real-world applications. Nevertheless, managing the trade-off between reward and safety during exploration presents a significant challenge. Improving reward performance through policy adjustments may adversely affect safety performance. In this study, we aim to address this conflicting relation by leveraging the theory of gradient manipulation. Initially, we analyze the conflict between reward and safety gradients. Subsequently, we tackle the balance between reward and safety optimization by proposing a soft switching policy optimization method, for which we provide convergence analysis. Based on our theoretical examination, we provide a safe RL framework to overcome the aforementioned challenge, and we develop a Safety-MuJoCo Benchmark to assess the performance of safe RL algorithms. Finally, we evaluate the effectiveness of our method on the Safety-MuJoCo Benchmark and a popular safe benchmark, Omnisafe. Experimental results demonstrate that our algorithms outperform several state-of-the-art baselines in terms of balancing reward and safety optimization.
@@ 1 Introduction
In this research, we aim to address the key question by leveraging the theory of gradient manipulation, wherein we conduct a detailed examination of the changes in gradients associated with reward and safety. Based on our theoretical analysis, we propose the projection constraint-rectified policy optimization (PCRPO) method, designed to alleviate the conflict between reward and safety optimization while maintaining a balance between their optimization levels, in which soft switching policy optimization through gradient manipulation is proposed and a slack technique is introduced for adjusting the emphasis on safety optimization. Particularly, we evaluate the effectiveness of our method on a multitude of challenging tasks, and conduct ablation experiments to thoroughly examine the performance of our method. The empirical findings suggest that our approach outperforms strong baselines concerning the balance between reward maximization and safety preservation.
@@ 3 Problem Formulation
Consequently, it is crucial to ascertain a method for delicately balancing the relationship between rewards and costs while simultaneously mitigating the adverse effects of conflicting gradients and minimizing the optimization oscillation on the overall performance of the approach. To satisfy the above requirements, a safe RL problem is rewritten as Equation (6), where $f_{rc_{i}}$ denotes the deviation between gradients, which could result in conflicting gradients and oscillations in the optimization process.
$\max_{\pi\in\Pi}$ $\ V^{\pi}_{r}(\rho),$ (6)
s.t. $V^{\pi}_{c_{i}}(\rho)\leq b_{i},\ \forall i=1,\ldots,n.$
$\min_{\pi\in\Pi}f_{rc_{i}}=f(\boldsymbol{\mathbf{g}_{r}},\boldsymbol{\mathbf{g}_{c_{i}}}).$
@@ 4 Method
In order to tackle the safe RL problem, as illustrated in Equation (6), it is imperative to first address the conflict between reward and cost gradients. To this end, we introduce a novel soft switching optimization solution that employs gradient manipulation to achieve a balanced relationship between these gradients. This approach incorporates a slack mechanism designed to gently optimize both reward and cost. Subsequently, we analyse the gradient change by soft switching. Then, the convergence analysis is provided. Lastly, we present a safe RL framework though gradient manipulation based on primal optimization, and a practical algorithm that effectively facilitates the implementation of our proposed method in real-world scenarios.
@@ 4.2 Gradient Analysis with Soft Switching
Furthermore, as illustrated in Figure 3 (c), the yellow dashed line signifies the updated gradient, $\mathbf{g}^{\prime}$, generated by Algorithm 1 of gradient surgery (Yu et al. 2020), the angle, $\theta^{s}$, between $\mathbf{g}^{\prime}$ and the original cost gradient, $\mathbf{g}_{c}$, remains greater than $90^{\circ}$. This observation implies that the relation between $\mathbf{g}^{\prime}$ and $\mathbf{g}_{c}$ continues to exhibit a conflicting nature, which may potentially result in inadequate handling of the original cost gradient. Based on the subsequent analysis as shown in Equation (17), our gradient manipulation approach demonstrates improved performance. It is important to note that Algorithm 1 of gradient surgery (Yu et al. 2020) does not consider cases where $\theta<90^{\circ}$, which might be insufficient for addressing optimization oscillations effectively. In the upcoming experiment section, we also provide ablation experiments to investigate the effectiveness of different gradient manipulation methods.
@@ 5.1 Experiments on Safety-MuJoCo Benchmark
We have developed a Safety-MuJoCo Benchmark [^1] https://github.com/SafeRL-Lab/Safety-MuJoCo.git based on MuJoCo (Todorov, Erez, and Tassa 2012) to evaluate the performance of safe RL algorithms. This benchmark differs from traditional safe RL benchmarks, such as Omnisafe [^2] https://github.com/PKU-Alignment/omnisafe.git (Ji et al. 2023) that is developed based on Safety Gym [^3] https://github.com/openai/safety-gym.git (Ray, Achiam, and Amodei 2019). In Omnisafe, the cost constraint is set as the velocity limit, and the reward is determined by the speed at which the robot runs. Conversely, our benchmark considers not only velocity constraints but also the health of the robot. For example, whether the robot falls or whether its joints exceed the limit values of motion control are taken into account. For a comprehensive description of these considerations, please refer to Appendix C. As depicted in Figure 4 (a) and (b), our method demonstrates remarkably superior performance in comparison to CRPO with respect to both reward maximization and safety preservation. Similarly, Figure 4 (c) and (d) illustrate our method’s significant improvement over CRPO in terms of reward and cost performance.
@@ 5.3 Ablation Experiments
As depicted in Figures 6 (c) and (d), we employ the gradient manipulation technique as outlined in Algorithm 1 of gradient surgery (Yu et al. 2020) for learning safety, named as SCRPO. The experimental results demonstrate that our proposed approach outperforms SCRPO in terms of safety and reward performance. These findings further corroborate the consistency of our theoretical analysis presented in the gradient analysis section.

### v3

@@ Abstract
Ensuring the safety of Reinforcement Learning (RL) is crucial for its deployment in real-world applications. Nevertheless, managing the trade-off between reward and safety during exploration presents a significant challenge. Improving reward performance through policy adjustments may adversely affect safety performance. In this study, we aim to address this conflicting relation by leveraging the theory of gradient manipulation. Initially, we analyze the conflict between reward and safety gradients. Subsequently, we tackle the balance between reward and safety optimization by proposing a soft switching policy optimization method, for which we provide convergence analysis. Based on our theoretical examination, we provide a safe RL framework to overcome the aforementioned challenge, and we develop a Safety-MuJoCo Benchmark to assess the performance of safe RL algorithms. Finally, we evaluate the effectiveness of our method on the Safety-MuJoCo Benchmark and a popular safe RL benchmark, Omnisafe. Experimental results demonstrate that our algorithms outperform several state-of-the-art baselines in terms of balancing reward and safety optimization.
@@ 1 Introduction
In this research, we aim to address the key question by leveraging the theory of gradient manipulation (Yu et al. 2020; Chen et al. 2021; Liu et al. 2021; Zhou et al. 2022), wherein we conduct a detailed examination of the changes in gradients associated with reward and safety. Based on our theoretical analysis, we propose the projection constraint-rectified policy optimization (PCRPO) method, designed to alleviate the conflict between reward and safety optimization while maintaining a balance between their optimization levels, in which soft switching policy optimization through gradient manipulation is proposed and a slack technique is introduced for adjusting the emphasis on safety optimization. Particularly, we evaluate the effectiveness of our method on a multitude of challenging tasks, and conduct ablation experiments to thoroughly examine the performance of our method. The empirical findings suggest that our approach outperforms strong baselines concerning the balance between reward maximization and safety preservation.
@@ 3 Problem Formulation
Consequently, it is crucial to ascertain a method for balancing the relation between rewards and costs while simultaneously mitigating the adverse effects of conflicting gradients and minimizing the optimization oscillation on the overall performance. To satisfy the above requirements, we optimize reward and safety performance while minimizing the deviation between reward and cost gradients. This helps prevent conflicting gradients and oscillations during the optimization process.
@@ 4 Method
In order to tackle the safe RL problem, as illustrated in Equation (4), it is imperative to first address the conflict between reward and cost gradients. To this end, we introduce a novel soft switching optimization solution that employs gradient manipulation to achieve a balanced relationship between these gradients. This approach incorporates a slack mechanism designed to smoothly optimize both reward and cost. Subsequently, we analyze the gradient change by soft switching. Then, the convergence analysis is provided. Lastly, we present a safe RL framework through gradient manipulation based on primal optimization, and a practical algorithm that effectively facilitates the implementation of our proposed method in real-world scenarios.
@@ 4.2 Gradient Analysis with Soft Switching
The gradient projection manipulation used in this study is inspired by gradient manipulation (Yu et al. 2020; Chen et al. 2021; Liu et al. 2021; Zhou et al. 2022), and we further leverage it for reward and safety balance. Specifically, as illustrated in Figure 3 (c), the yellow dashed line denotes the updated gradient, $\mathbf{g}^{\prime}$, generated by Algorithm 1 of gradient surgery (Yu et al. 2020), the angle, $\theta^{s}$, between $\mathbf{g}^{\prime}$ and the original cost gradient, $\mathbf{g}_{c}$, remains greater than $90^{\circ}$. This observation implies that the relation between $\mathbf{g}^{\prime}$ and $\mathbf{g}_{c}$ continues to exhibit a conflicting nature, potentially resulting in inadequate handling of the original cost gradient. Based on the subsequent analysis as shown in Equation (16), our gradient manipulation approach demonstrates improved performance. Algorithm 1 of gradient surgery (Yu et al. 2020) does not consider cases where $\theta<90^{\circ}$, which might be insufficient for addressing optimization oscillations effectively. Thus, on the basis of gradient manipulation (Yu et al. 2020; Chen et al. 2021; Liu et al. 2021; Zhou et al. 2022), we consider both conflicting and non-conflicting gradient scenarios in safe RL. In the upcoming experiment section, we also provide ablation experiments to investigate the effectiveness of gradient manipulation methods regarding reward and safety optimization.
@@ 5.1 Experiments on Safety-MuJoCo Benchmark
We have developed a Safety-MuJoCo Benchmark [^1] https://github.com/SafeRL-Lab/Safety-MuJoCo.git based on MuJoCo (Todorov, Erez, and Tassa 2012; Towers et al. 2023) to evaluate the performance of safe RL algorithms. This benchmark differs from traditional safe RL benchmarks, such as Omnisafe [^2] https://github.com/PKU-Alignment/omnisafe.git (Ji et al. 2023) that is developed based on Safety Gym [^3] https://github.com/openai/safety-gym.git (Ray, Achiam, and Amodei 2019). In Omnisafe, the cost constraint is set as the velocity limit, and the reward is determined by the speed at which the robot runs. Conversely, our benchmark considers not only velocity constraints but also the health of the robot. For example, whether the robot falls or whether its joints exceed the limit values of motion control are taken into account. For a comprehensive description of these considerations, please refer to Appendix C. As depicted in Figure 4 (a) and (b), our method demonstrates remarkably superior performance in comparison to CRPO with respect to both reward maximization and safety preservation. Similarly, Figure 4 (c) and (d) illustrate our method’s significant improvement over CRPO in terms of reward and cost performance.
@@ 5.3 Ablation Experiments
As depicted in Figures 6 (c) and (d), we employ the gradient manipulation technique as outlined in Algorithm 1 of gradient surgery (Yu et al. 2020) for learning safety, named as SCRPO. The experimental results demonstrate that PCRPO outperforms SCRPO in terms of safety and reward performance. These findings further corroborate the consistency of our theoretical analysis presented in the gradient analysis section.